OpenStudy (barrelracing):

Which of the following is a solution of x2 + 2x + 4? a) -2 + 2i squareroot 3 b) 2 +2i squareroot 3 c) 1 - I squareroot 3 d) -1 + I squareroot 3

Nnesha (nnesha):

you need to use quadratic formula \[\huge\rm x =\frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\] a ,b,and c values are \[\large\rm Ax^2+Bx+C=0\] a=leading coefficient b= middle term c= constant term

OpenStudy (barrelracing):

\[x=\frac{ -2x \pm -12 }{ 8}\]

Nnesha (nnesha):

middle term is b which is 2 we don't the x variable and what is a= ?

Nnesha (nnesha):

remember c is constant term and a is leading coefficient \[\huge\rm \color{reD}{A}x^2+\color{blue}{B}x+\color{salmon}{C}=0\] \[\huge\rm \color{ReD}{1}x^2+\color{blue}{2}x+\color{salmon}{4}=0\]